### v1

# Moran Mizrahi1 Chen Shani2 Gabriel Stanovsky1 Dan Jurafsky2 Dafna Shahaf1 1

Abstract
Large Language Models (LLMs) excel at countless tasks, yet struggle with creativity. In this paper, we introduce a novel approach that couples LLMs with structured representations and cognitively inspired manipulations to generate more creative and diverse ideas. Our notion of creativity goes beyond superficial token-level variations; rather, we explicitly recombine structured representations of existing ideas, allowing our algorithm to effectively explore the more abstract landscape of ideas. We demonstrate our approach in the culinary domain with $\mathcal{D}$ishCover, a model that generates creative recipes. Experiments comparing our model’s results to those of GPT-4o show greater diversity. Domain expert evaluations reveal that our outputs, which are mostly coherent and feasible culinary creations, significantly surpass GPT-4o in terms of novelty, thus outperforming it in creative generation. We hope our work inspires further research into structured creativity in AI.
[Uncaptioned image] Cooking Up Creativity: A Cognitively-Inspired Approach for Enhancing LLM Creativity through Structured Representations
Moran Mizrahi1 Chen Shani2 Gabriel Stanovsky1 Dan Jurafsky2 Dafna Shahaf1 1Hebrew University of Jerusalem  2Stanford University {moranmiz, dshahaf, gabriel.stanovsky}@cs.huji.ac.il  {cshani, jurafsky}@stanford.edu
1 Introduction
Figure 1: The $\mathcal{D}$ishCover pipeline for creative recipe generation (with LLM-based components shaded in blue) takes as input two idea descriptions. Each description is mapped to a set of specific recipes (§ 4.1), which are parsed into tree representations (§ 4.2). These trees are subsequently combined using a minimal edit distance algorithm (§ 4.3), assessed for value, and ranked based on novelty scores (§ 4.4). Finally, the highest-ranked trees are translated back into natural language recipes (§ 4.5).
Large Language Models (LLMs) excel at generating fluent, coherent text and performing tasks that draw on extensive world knowledge. However, they often struggle to generate truly creative ideas Franceschelli and Musolesi (2024); Chakrabarty et al. (2024); Tian et al. (2024b); Zhao et al. (2024).
In creativity research, creative outputs are typically defined as those that are both novel (unexpected and original) and valuable (useful, relevant, or effective) Mumford (2003); Boden (2004). However, due to their reliance on vast repositories of existing data, LLMs inherently follow learned patterns, making them prone to producing repetitive or predictable outputs that lack genuine novelty. Ironically, attempts to explicitly instruct LLMs to “think more creatively” often leads them to generate invalid or hallucinated solutions that could mislead uninformed users Wang et al. (2024a); Jiang et al. (2024). Together, these limitations make creative generation a persistent challenge for LLMs.
The temperature parameter of LLMs controls the amount of randomness, and is often claimed to be the creativity parameter. However, creativity encompasses much more than mere randomness; a recent study Peeperkorn et al. (2024) found that while higher temperatures weakly correlate with increased novelty, their actual influence on overall creativity remains subtle and limited.
Much recent work has shown that combining structured knowledge (such as knowledge graphs) with LLMs can significantly improve their performance Pan et al. (2024); Wang et al. (2024b); Feng et al. (2023); Sun et al. (2023). In particular, a line of works uses an LLM to parse text into structured representation, manipulates this representation, and (optionally) applies the LLM again to translate the result into natural language, especially in inference and reasoning tasks Yang et al. (2023); Besta et al. (2024); Zelikman et al. (2023); Zhang et al. (2025).
In this work, we show that surprisingly, incorporating structure can also improve the LLM’s creativity and diversity. We stress that we do not mean creativity and diversity on the lexical (token) level; Rather, we want the model to be creative on a more abstract level, within the realm of concepts (or the “landscape of ideas”, so to speak).
Our approach is illustrated in Figure 1. Similar to parsing-based approaches Tian et al. (2024a); Zhao et al. (2023); Li et al. (2024), we start by deriving structured representations from textual inputs. Next, we manipulate these structured representations to reach creative parts of the idea space. We draw inspiration from the human creative process, and focus on recombination – a fundamental principle in creativity research, which posits that novel ideas often emerge by merging existing concepts in unexpected ways Guilford (1967); Utterback (1996); Ahuja and Morris Lampert (2001). For example, combining pizza preparation method with the flavors of alfajores cookies might yield a brand-new “alfajores pizza”, and combining a sofa with a bookshelf might result in multifunctional furniture. We sample ideas from the space of recombinations, and evaluate them based on novelty and value Finke et al. (1996); Sawyer and Henriksen (2024).
We demonstrate our paradigm in the culinary domain, introducing $\mathcal{D}$ishCover, a model for creative generation of recipes. Figure 2 presents examples of recombinations generated by it.
Beyond the immediate scope of cooking, we believe this paradigm holds promise for extending creative and diverse generation to a wide range of domains, from product design, narrative construction, and scientific discovery to artistic composition. In summary, our contributions are:
  • We introduce a novel paradigm to enhance LLM creativity by extracting structured representations from natural language and manipulating them (inspired by cognitive processes), going beyond mere token-level variability.
  • We demonstrate our approach in the culinary domain with $\mathcal{D}$ishCover, a model that recombines recipes to generate creative ones.
  • We curate a 5K-recipe dataset generated by $\mathcal{D}$ishCover, providing a valuable resource for future work on creative generation. We make both the code and data publicly available. [^1] https://github.com/moranmiz/Cooking-Up-Creativity
  • Through systematic experiments, we show that $\mathcal{D}$ishCover’s generations are significantly more diverse compared to baseline SOTA LLM outputs. Most recipes generated by both models are deemed valuable (appropriate and coherent), although the baseline achieves better scores on an open-ended task. Most importantly, our outputs significantly surpass the baseline in terms of novelty, resulting in more creative culinary ideas. These findings are supported by both automated metrics and domain expert evaluations.
Figure 2: Examples of new recipe ideas generated by $\mathcal{D}$ishCover. Each example consists of a pair of input dishes and their best corresponding generated recipe idea. The generated recipes are presented as concise summaries to conserve space. Images were generated using OpenAI’s DALL·E.
2 Background: Human Creativity
The field of human creativity has been extensively studied, identifying numerous principles that drive innovation. In designing our model, we relied on the following principles:
Generation & Evaluation. A common yet effective model of creative thinking is the two-stage process generation & evaluation, which suggests that creativity begins with divergent thinking (free idea generation), followed by convergent thinking, where the most promising ideas are selected and refined Finke et al. (1996); Sawyer and Henriksen (2024). We incorporate this as the conceptual backbone of our model, implementing a generative component that produces a broad set of ideas, followed by an evaluative component that identifies those with the greatest creative potential.
Recombination of Ideas. We base our work on a prominent idea-generation method: recombination, where elements from existing ideas are merged to create novel concepts Guilford (1967); Koestler (1964). Our model strategically recombines elements from pairs of existing seed ideas to spark unexpected connections.
Creativity Assessment: Novelty & Value. After generating many ideas, the challenge is determining which are genuinely creative. Numerous studies have examined the complexities of assessing creativity in both humans and computational systems Said-Metwaly et al. (2017); Lamb et al. (2018). A widely accepted definition of creativity frames it as the intersection of novelty and value Mumford (2003); Boden (2004, 2009); Lamb et al. (2018). Novelty ensures that an idea is surprising or unconventional, while value signifies it is useful in its intended context.
Measuring Novelty and Value Automatically.
Novelty can be assessed by identifying how uncommon an idea is within a dataset Heinen and Johnson (2018); Kenett (2019); Doboli et al. (2020). Evaluating value, however, is highly domain-dependent, often considered the “holy grail” of computational creativity Boden (2004); Ritchie (2007); Jordanous (2012). Thus, we consider value assessment as a domain-specific task.
3 Problem Definition
Innovation often involves combining existing ideas to create novel ones. This process, often referred to as “conceptual blending” or “creative recombination,” is central to innovation, and the focus of our work. We now introduce key elements of our formulation.
Given a domain where ideas can be expressed in a structured format (e.g., cooking recipes, instruction manuals, computer programs), let $\mathcal{I}$ denote the theoretical set of all possible ideas within that domain – both existing and yet-to-be-discovered. $\mathcal{I}$ represents the entire conceptual space of ideas that adhere to the domain’s structural and logical constraints, encompassing all valid possibilities. In addition, let $I\subset\mathcal{I}$ be a set of ideas that have been recorded or are known within the field.
Definition 1 (Recombination Function).
Recombination function $\mathcal{C}$ takes as input two structured ideas $i_{a},i_{b}\in I$ and produces a set of new combinations $I_{ab}\subseteq\mathcal{I}$ such that each $i\in I_{ab}$ is a different mixture of $i_{a},i_{b}$.
The exact definition of “mixture” depends on the representation. For example, when we transition from representation $i_{a}$ to representation $i_{b}$ with a minimal edit distance procedure, the intermediate steps can be viewed as mixtures of $i_{a}$ and $i_{b}$, blending elements of both in varying proportions as we move through the transformation.
Definition 2 (Evaluation Function).
The output of a recombination is a set of potential innovations $I_{ab}$, which can be evaluated with evaluation function $E:\mathcal{I}\rightarrow\mathbb{R}$. The innovation can be evaluated based on criteria such as novelty and utility.
Definition 3 (Retrieval of Ideas from Descriptions).
Ideas are often expressed in different levels of abstraction and granularity. Let $m$ be a function that matches an idea description $d$ to relevant known ideas from $I$, $m(d)\subseteq I$. For example, $m$ could match the textual description “lasagna” to all lasagna recipes.
The formal optimization problem can thus be stated as: Given two idea descriptions $d_{a},d_{b}$, find
$\operatorname*{argmax}_{i\in\mathcal{C}(i_{a},i_{b})\ \mid\ i_{a}\in m(d_{a}),i_{b}\in m(d_{b})}E(i)$
Figure 2 illustrates examples of generated ideas in the domain of cooking recipes, along with the idea descriptions used to create them. For example, combining broccoli salad recipes and burrito recipes resulted in a recipe for a tortilla filled with cheese, broccoli, strawberries and pistachios.
4 Model
In this section, we introduce $\mathcal{D}$ishCover, our model for automatically generating innovative recipes.1 Figure 1 illustrates our methodology. The input consists of two seed inspirations (idea descriptions $d_{a}$, $d_{b}$) [^2] Note that more inspirations can be used if desired.. Each idea description is mapped to a set of specific recipes (instances of the idea, for example different lasagna recipes; § 4.1).
These recipes are first translated into tree representations with the help of an LLM (step (I) in Figure 1, § 4.2). To generate new ideas, we recombine these trees to produce new candidate ideas with a minimal edit-distance algorithm (step (II), § 4.3). Then, utilizing the Novelty & Value principle, we refine the candidate ideas to assess their value and rank them based on their novelty scores (step (III), § 4.4). Finally, the highest-ranked trees are translated back into natural language recipes and refined using an LLM (step (IV), § 4.5). We now provide more details about each step.
4.1 Sampling Seed Ideas (Step I)
We selected the 100 most popular dishes (e.g., chicken salad, cheesecake) that span different categories (e.g., appetizers, desserts, main courses) from the Recipe1M+ dataset Marın et al. (2021). The average number of appearances in the dataset is 2,576.33 recipes per dish.
To keep the financial costs of using an LLM manageable in the next stage of our model, we sampled 30 recipes per dish, resulting in a total of 3K recipe samples. To ensure both diversity and representativeness, we selected 15 recipes at random to capture the typical version of each dish, and 15 more to maximize diversity. Implementation of diversification can depend on the representation and the domain; in our case, we use the GMM algorithm Ravi et al. (1994) over recipe embeddings from a Sentence-BERT model fine-tuned on recipe data (see details in Appendix A).
Figure 3: Tree representations of lasagna and chocolate pie recipes. Analogous parts are highlighted, showing structural similarities that the minimal edit distance algorithm is more likely to preserve when transforming one tree into another.
4.2 Text to Tree (Step II)
Cooking recipes, like science experiments, assembly manuals, and game instructions, are procedural texts. These texts typically consist of a sequence of steps accompanied by the objects needed to perform them. A common way to represent procedural texts is with a tree Jermsurawong and Habash (2015); Maeta et al. (2015), where leaf nodes correspond to needed objects (ingredients, in our case), and the internal nodes represent the actions performed on them. Figure 3 shows simple lasagna and chocolate pie recipes in this tree format.
To parse recipe text into tree, we prompted GPT-4o, leveraging its extensive world knowledge and code generation capabilities with a chain-of-thought approach. See Appendix BC for full details and corresponding prompts. The total cost of generating tree representations for 3K recipes was approximately \$40.
An initial pass revealed that 1,347 (44.9%) of the resulting trees were invalid due to issues such as orphan nodes, multiple outgoing edges from a single node, or incorrect edge directions. To address this, we implemented a correction step in which we removed problematic edges and instructed the model to reconsider them. This improved validity to 95% (2,850 trees). We then evaluated the final trees using 50 random recipes. A gold-standard tree was created for each recipe, and we automatically compared the predicted trees with the gold trees in terms of node and edge matching. For nodes, we achieved a Precision of 0.985, Recall of 0.956, and F1-Score of 0.969. For edges, we obtained a Precision of 0.951, Recall of 0.909, and F1-Score of 0.93. Overall, these results demonstrate the effectiveness of our approach in translating procedural texts into a structured tree representation, making them suitable for further manipulation and analysis.
Figure 4: Illustration of our tree-based edit-distance approach for generating new recipe ideas by transforming a simple lasagna tree (left) into a simple chocolate pie tree (right). Intermediate “merged” trees are produced for each edit step, yielding novel dishes such as a basil chocolate lasagna or a chocolate lasagna encased in a pie crust. Recipe images were generated using OpenAI’s DALL·E model.
4.3 Generate Ideas (Step III)
In this section we operationalize the recombinaction function $\mathcal{C}$. We generate novel recipes by blending recipe trees with a minimal edit-distance algorithm. The key idea behind this method is that by examining the step-by-step transformation between two concepts, we can discover intermediate forms that blend features of both.
In the case of recipe generation, we employ the Zhang–Shasha algorithm, which computes the minimal edit distance between trees Zhang and Shasha (1989); Bille (2005). Given two recipe trees $i_{a}$ and $i_{b}$, we compute their minimal edit distance and record all operations required to transform $i_{a}$ into $i_{b}$. Each operation sequence produces intermediate trees that represent novel “merged” ideas, from which we randomly select one as our new recipe. Figure 4 illustrates an example sequence transforming a simple lasagna tree into a simple chocolate pie tree. One intermediate variant might be a chocolate lasagna with basil; another could feature a chocolate lasagna encased within a crust.
A key advantage of the minimal edit distance approach is its ability to preserve the structural roles of ingredients and cooking steps. For example, in Figure 3, both the lasagna and chocolate pie recipes include a “topping” action (marked in color). Using a minimal edit distance ensures that inserting whipped cream aside of tomato sauce and cheese is more likely, as placing whipped cream elsewhere would increase the overall edit cost. See implementation details in Appendix D.
Note that stopping at different points in the transformation process can create unique dishes (see Figure 4). Additionally, shuffling the order of edits can generate entirely new intermediate ideas.
4.4 Evaluate Ideas (Step IV)
Now that we have a set of candidate innovations created through reformulation, we evaluate each generated recipe. As noted in Section 3, the evaluation function $E$ should take into account novelty and value. Specifically, we chose to view value as a constraint, and novelty as the optimization objective; i.e., we wish to rank by novelty all candidates that pass a value threshold (i.e., make sense).
In the culinary domain, a recipe is often considered creative if it introduces unexpected ingredient or technique combinations (novelty) while still producing a delicious, coherent dish (value, utility). In the following, we operationalize these criteria. Our implementation focuses on recipes, but we believe the principle can be generalized to other domains.
4.4.1 Value Constraint
To assess value (taste) of a recipe, we follow Varshney et al. (2019), and check if its ingredients pair well with each other. We use flavorDB dataset Garg et al. (2018), which catalogs taste molecules for raw ingredients. According to this work, two ingredients pair well if they share a larger proportion of taste molecules; we compute a pairing score for each pair of raw ingredients. Since this dataset does not cover processed ingredients, we also use FoodData Central Fukagawa et al. (2022) to infer their raw components. We define the pairing score between two composite ingredients as the lowest score among their constituent raw-ingredient pairs. After exploration, we chose 0.3 as our value threshold. If the score falls below this threshold, we consider the pairing problematic. We iteratively remove the ingredient with the highest number of low-scoring pairings until no further collisions remain (alternatively, one might choose to remove these candidates altogether). See details in Appendix E.
4.4.2 Ranking by Novelty
Our intuition is that novel recipes include ingredients and actions that do not often appear together. Thus, we formulate a measure of surprise, inspired by the inverse document frequency (idf) concept in tf-idf Ramos et al. (2003).
Given recipe tree $T$ with nodes $E_{T}$ (ingredients and actions), let $N_{e}$ be the number of recipes in the repository that include element $e$. For each element $e^{\prime}\in E_{T}\setminus e$ , $df_{e}(e^{\prime})$ denotes the number of recipes containing both $e,e^{\prime}$. We then define [^3] To prevent score inflation due to typos or extremely rare ingredients, we exclude any ingredient that appears exceptionally infrequently in the entire dataset.:
$idf_{e}(e\textquoteright)=log\Bigl{(}\frac{N_{e}}{df_{e}(e^{\prime})}\Bigr{)}$
A higher $idf_{e}(e\textquoteright)$ ’ ) score indicates that $e^{\prime}$ is more unique relative to recipes containing $e$.
To compute novelty of an element in $T$, we compute all $idf_{e}(e\textquoteright)$ ’ ) scores and sum the top ten. To compute the novelty of $T$ itself, we sum the top-10 element novelties. For example, in Figure 2, the chicken and mango delight recipe was ranked as novel due to ingredient combinations like mango, crescent roll, nuts and broccoli, and actions such as steaming, blanching and unrolling.
4.5 Tree to Text (Step V)
After identifying the highest-ranked trees, we convert them back into natural-language recipes using an LLM. Similar to Section 4.2, we employ a chain-of-thought approach, instructing GPT-4o to translate the tree (encoded in DOT format) into a structured recipe, including a title, ingredient list, and step-by-step instructions. We then prompt the model to refine and correct the text for coherence, quantity adjustments, ingredient consistency, and overall readability (see Appendix F for full prompts). This approach required a total cost of approximately \$42 to process 1000 recipe trees.
5 Creative Recipe Dataset
We used our model to generate a new dataset of recipes. We first identified 100 popular dishes spanning different categories. For each dish, we sampled 30 recipes and converted them into tree structures. Next, we sampled 1,000 dish pairs, ensuring that pairs include dishes from different categories. For each pair of dishes, $d_{a}$ and $d_{b}$, every recipe pair in $m(d_{a})\times m(d_{b})$ was used to generate six blended trees, producing up to 5,600 trees per dish pair ($\sim$5.5M in total). We selected the five highest-ranked trees from each set and converted them back into natural language recipes, resulting in a dataset of 5K recipes.
6 Evaluation
We now turn to evaluate $\mathcal{D}$ishCover by investigating the following research question: How do the recipes $\boldsymbol{\mathcal{D}}$ishCover generated compare to those generated by a SOTA LLM (GPT-4o)?
To answer this, we examine two key aspects. First, how the outputs compare in terms of diversity. Specifically, we investigate whether our approach mitigates the well-documented issue of repetitiveness in LLMs, which is a critical challenge in creative generation. More importantly, we assess how our outputs compare those of GPT-4o in terms of creativity. Creativity requires outputs that are both valuable (make sense) and novel (unexpected). We compare our model’s outputs with those of GPT-4o [^4] Version: gpt-4o-2024-08-06 (latest stable version of the model at the time of writing this paper). on two tasks:
Experiment 1. We evaluated GPT-4o and $\mathcal{D}$ishCover on their ability to generate creative recipes combining a given dish pair. The evaluation included randomly selected 10 dish pairs. For each pair, we selected 5 recipes generated by GPT-4o and compared them to the top 5 recipes generated by $\mathcal{D}$ishCover, resulting in a total of 50 recipes per model.
Experiment 2. To broaden the scope of our analysis, we wish to evaluate the most creative recipes GPT-4o and $\mathcal{D}$ishCover could generate in general, without limiting them to a given input pair. We used 100 recipes from each model. We instructed GPT-4o to generate 100 different creative recipes in a single chat session. For $\mathcal{D}$ishCover, we selected 100 recipes from our 5K-recipe dataset, which includes a novelty score for each generated recipe. To ensure that the dishes do not come from a very small number of inputs, we employed simulated annealing Bertsimas and Tsitsiklis (1993) to maximize recipe novelty while enforcing a constraint on the maximum number of appearances per dish.
Note that Experiment 1 is a more unconventional task, one that the model is less likely to have encountered during training. We evaluate the outputs of both experiments using qualitative and automated analyses as well as human annotations.
6.1 Experimental Details
Here we describe the process of selecting prompting parameters for GPT-4o as well as the human evaluation setup, both of which were consistently applied across both experiments.
Prompt and Temperature Selection. LLMs are known to be sensitive to prompt paraphrases Sclar et al. (2023); Mizrahi et al. (2024); Voronov et al. (2024). To optimize our parameters, we first determined the temperature value by generating outputs from three random prompts. After testing various settings, we selected $t=1$ as the highest temperature that produced coherent recipes, noting that higher values resulted in gibberish. With the temperature fixed, and following the methodology of Mizrahi et al. (2024), a team member curated 100 prompts per experiment, generating three recipes for each prompt. The 300 resulting recipes were then reviewed, and the more unique outputs were tagged. Ultimately, the final prompt was chosen for its ability to consistently generate creative recipes across all outputs (see the chosen prompts for both experiments in Appendix G).
Human Evaluation Setup. We used Prolific to recruit and manage a total of 48 participants for both experiments. We pre-screened participants based on their cooking frequency, comfort with adjusting recipes, and ability to judge creative outcomes. Participants were paid an estimated hourly rate of £9, adhering to ethical compensation standards. Participants were asked to rate random recipes drawn from the two models on various value and novelty aspects (see the list of question in Appendix H). To reduce cognitive load, they viewed short recipe summaries (also generated using GPT-4o), with the option to view full recipes if desired. To mitigate ranking noise, each recipe was rated by five annotators, and final scores were determined using median ranks. The novelty and value scores were set to be the means of their corresponding questions, and the value score was also binarized using a threshold of 4 on three related value questions. On average, each participant rated 31.25 recipes (std=33.827). To ensure quality, we excluded participants who, on average, completed reading a recipe and answering its eight questions in less than 45 seconds.
Figure 5: Ingredient frequencies in the original recipe repository (in pink) and the models’ generated recipes. The black histogram shows $\mathcal{D}$ishCover closely follows the repository’s distribution. GPT-4o (blue) shows spikes, highlighting bias to certain ingredients.
6.2 Diversity in Generated Recipes
Diversity plays a pivotal role in creative generation, reflecting how broadly and flexibly new recipes adapt and transform original concepts. In this part of our evaluation, we use both qualitative and automated analyses to compare $\mathcal{D}$ishCover’s outputs with those generated by GPT-4o, examining how each model integrates concepts from different dishes and their level of repetitiveness.
Deep vs. Shallow Merging of Recipe Concepts.
Both $\mathcal{D}$ishCover and GPT-4o demonstrated the ability to merge ideas from different dishes, but we noticed their outputs exhibited different types of integration. $\mathcal{D}$ishCover consistently produced more cohesive integrations, where ingredients from both source dishes were woven together into a single, unified recipe. In contrast, GPT-4o tended toward “shallow merges”, typically preparing each dish separately before combining them at the end. For example, when asked to create a hybrid of muffins and orange salad, GPT-4o invariably proposed baking muffins, preparing an orange salad, and then serving them together, placing the salad on top or on the side. In contrast, $\mathcal{D}$ishCover generated recipes like a mixed vegetable bake incorporating elements from both dishes, a citrus cake with an orange-based sauce, and a salad that integrated muffin ingredients.
This difference was also reflected in the average ingredient counts per recipe: GPT-4o’s recipes had nearly twice as many ingredients as $\mathcal{D}$ishCover’s: in the first experiment, $\mathcal{D}$ishCover used 12.3 ingredients on average (std=3.71) compared to GPT-4o’s 24.98 (std=5.2); in the second experiment, $\mathcal{D}$ishCover used 13.32 ingredients on average (std=4.39) vs. GPT-4o’s 23.79 (std=3.25). For reference, the average ingredient count in the recipe repository was lower, 9.33 (std=4.31). This imbalance suggests that GPT-4o tends to generate separate sub-dishes, each with its own ingredients, and then combine them into a final dish rather than truly blending the culinary ideas.
Figure 6: t-SNE visualization for recipe embeddings, experiment 1. Shapes denote dish pairs, colors denote models. The blue clusters (GPT-4o) of the same shape appear tight and localized, while the pink ($\mathcal{D}$ishCover) are scattered, showing higher diversity across $\mathcal{D}$ishCover’s outputs.
Figure 7: t-SNE visualization for recipe embeddings, experiment 2. Gray points represent 1K random recipes from the general repository. Again, $\mathcal{D}$ishCover’s outputs are more diverse.
Fixation on Structures and Ingredients in GPT-4o’s Outputs.
GPT-4o showed a strong fixation on certain structures and ingredients. When asked to merge two dishes, it often followed the same approach in multiple attempts. For instance, when asked to fuse “lentil soup” and “jam”, GPT-4o repeatedly generated variations of lentil soup served alongside a separately prepared jam, typically plated in the same way (e.g., placing a dollop of jam in the center of the soup). This fixation persisted when asked to combine specific given recipes for each dish, and even when asked to merge broader dish types (e.g., a general soup with a general dip).
Similarly, GPT-4o repeatedly used the same (uncommon) ingredients. For instance, 56% of its recipes included smoked paprika (compared to just 0.375% in the repository), 47% used mint, 41% maple syrup, and 29% pomegranate molasses. Figure 5 shows ingredient frequencies in GPT-4o and $\mathcal{D}$ishCover recipes against the repository baseline. While $\mathcal{D}$ishCover’s distribution aligns closely with the original data, GPT-4o has a strong bias toward certain low-frequency ingredients.
Quantifying Diversity via Tree Distances.
We computed the average normalized tree edit distance Rico-Juan and Micó (2003) between generated recipes after converting them into hierarchical tree structures. In both experiments, GPT-4o’s outputs had significantly lower tree distances than $\mathcal{D}$ishCover’s, indicating higher similarity. In the first experiment, we calculated the average edit distance for each dish pair separately, finding that $\mathcal{D}$ishCover had an average tree distance of 132.14, while GPT-4o’s was 89.35 (p-value=3.6e-05, paired t-test). In the second experiment, across all outputs, $\mathcal{D}$ishCover again exhibited greater diversity, with an average tree distance of 140.25 compared to GPT-4o’s 129.55 (p-value<1e-50, two-sample t-test).
Quantifying Diversity via Embeddings.
We also analyzed the embeddings of the generated recipes using a fine-tuned Sentence-BERT model specialized for cooking recipes (see Appendix A). Figure 6 presents a t-SNE visualization for the first experiment, where each shape represents a dish pair, and color indicates the model ($\mathcal{D}$ishCover in pink, GPT-4o in blue). GPT-4o’s recipes formed tight clusters for each pair, whereas $\mathcal{D}$ishCover’s were more dispersed. A similar pattern emerges in experiment 2 (Figure 7). By construction, $\mathcal{D}$ishCover’s embeddings are widely distributed. The interesting thing to notice, however, is that GPT-4o’s embeddings tend to cluster closely. Figure 8 reinforces these findings with heatmaps of cosine similarities among all recipe embeddings. In the first experiment, $\mathcal{D}$ishCover’s outputs had an average similarity of 0.387 (std=0.120), while GPT-4o’s were higher at 0.659 (std=0.121). This pattern held in the second experiment ($\mathcal{D}$ishCover: 0.402, std=0.110; GPT-4o: 0.731, std=0.078).
Figure 8: Four heatmaps illustrating pairwise cosine similarities of recipe embeddings for each model across both experiments. GPT-4o’s heatmaps (right) display more areas of high similarity, indicating $\mathcal{D}$ishCover’s higher diversity.
6.3 Human annotation
While diversity is crucial, genuine creativity in recipe generation also depends on value and novelty. To assess these aspects, we built on the cooking expert annotations from the human evaluation.
In terms of value, both models produced outputs that were largely deemed valuable. In the first experiment, 80% of $\mathcal{D}$ishCover’s recipes were classified as valuable, with an average value score of 4.320, compared to 82% for GPT-4o (4.326). In the second experiment, 85% of $\mathcal{D}$ishCover’s outputs were classified as valuable (4.36), compared to 98% for GPT-4o (4.51). This makes sense, as we noted that the second experiment (open-ended prompts) is more similar to the data GPT-4o has encountered during training, and also that it tends to output relatively similar recipes.
For novelty, we considered only the recipes that were classified as valuable; this reflects a use-case of a cook skimming a list of recommended recipes, quickly judging their sensibility, delving only into those that have potential. In both experiments, $\mathcal{D}$ishCover had significantly outperformed GPT-4o in this regard. In the first experiment, $\mathcal{D}$ishCover’s average novelty score was 3.53, compared to GPT-4o’s 3.146 (p-value=0.0009). In the second experiment, $\mathcal{D}$ishCover’s average novelty score was 3.612, compared to GPT-4o’s 3.141 (p-value=9.2E-08).
Examining the second experiment’s valuable outputs further, $\mathcal{D}$ishCover dominated the top quartile of novelty scores, making up 75.55% of the highest-rated recipes, while GPT-4o prevailed in the lower two quartiles, accounting for 71.74% of the less novel outputs. Moreover, among the 37 valuable recipes with a novelty score of 4 or higher, $\mathcal{D}$ishCover contributed 32, leaving only five from GPT-4o. Figure 2 showcases five $\mathcal{D}$ishCover recipes from this high-novelty set. These results strongly suggest that while both models produce mostly valuable recipes (GPT-4o more so, in the open-ended case), $\mathcal{D}$ishCover has a clear advantage in generating truly creative ones.
7 Related Work
Our approach builds on recent parsing-based methods that guide LLMs to map natural language into structured forms, boosting LLM performance in tasks such as constituency parsing Tian et al. (2024a) and information extraction Zhao et al. (2023); Li et al. (2024). Similar to research integrating LLMs with Knowledge Graphs (KGs) to improve inference and reasoning Wang et al. (2024b); Feng et al. (2023); Jiang et al. (2023); Sun et al. (2023), we incorporate domain-specific knowledge into structured representations to provide LLMs with clearer, context-rich signals. Our work also aligns with models that parse text into structured knowledge, manipulate these representations, and (optionally) convert them into natural language to enhance LLM capabilities Yang et al. (2023); Besta et al. (2024); Zelikman et al. (2023); Zhang et al. (2025). We surprisingly show that similar techniques can augment creativity and diversity.
Our focus on increasing LLM output diversity complements efforts that promote variation in generation via human feedback Chung et al. (2023), in-context learning Zhang et al. (2024), or KG-based interventions Liu et al. (2021); Hwang et al. (2023b); Liu et al. (2022). Additionally, our work aligns with broader research aimed at fostering creativity in LLMs. One line of work leverages hallucinations for new ideas Jiang et al. (2024); Yuan and Färber (2025). Another one draws on insights from human creativity research, incorporating techniques such as constraints Lu et al. (2024), associative thinking Mehrotra et al. (2024), role-playing Chen et al. (2024), and brainstorming Summers-Stay et al. (2023); Chang and Li (2025); Rana and Cheok (2025). Similarly, we focus on recombination.
Beyond improving LLMs, researchers have also explored their use as creative aids for writers Mirowski et al. (2023); Yuan et al. (2022); Chakrabarty et al. (2023); Wan et al. (2024), visual artists Ko et al. (2023), and even humorists Wu et al. (2025). However, while these tools boost users’ sense of creativity, their limited diversity may homogenize the ideas produced across different individuals Anderson et al. (2024).
Recent work has explored the use of LLMs in the culinary domain Ma et al. (2024); Zhou et al. (2024); Ataguba and Orji (2025); Hwang et al. (2023a), including several early efforts in recipe generation H. Lee et al. (2020); Antô et al. (2020); Bień et al. (2020). Prior to the advent of LLMs, computational recipe generation focused primarily on proposing novel ingredient combinations, often neglecting complete cooking instruction generation Morris et al. (2012); Amorim et al. (2017); Cromwell et al. (2015); Varshney et al. (2019).
8 Discussion and Conclusions
We introduce a novel paradigm for enhancing creative generation in LLMs using structured recombination at the idea level rather than token-level.
Implications for AI Creativity and Structured Knowledge. A core insight from our work is that structured representations provide a mechanism to introduce meaningful variability at a higher level of abstraction. Attempts to enhance creativity in LLMs often focus on increasing token-randomness (e.g., through temperature), which is not sufficient for creativity Peeperkorn et al. (2024). We focus on recombination, a well-established principle in creativity research. By operationalizing this principle computationally, we show that incorporating structured knowledge enables a model to navigate the space of ideas more effectively, leading to novel and valuable outputs.
Implications for Sampling. Another promising application of our approach is in structured sampling and search-based generation tasks. Sampling methods in LLMs often rely on stochastic techniques that operate at the token level (e.g., nucleus or top-k sampling) to introduce variability. In contrast, we diversify by sampling over abstract representations rather than raw tokens. We introduce a controlled yet flexible way of navigating idea spaces. This approach to sampling can enhance a range of applications that require balancing novelty and coherence.
Generalizability to Other Domains. Beyond cooking, our framework can generalize to any domain where structured knowledge can guide creative recombination, such as product design, scientific discovery, or narrative construction. Tree representations apply naturally to many of these domains, and our approach can also be extended to other structures, such as sequences, general graphs, or other formal structures. In addition to idea generation, our framework can support synthetic data simulation and augmentation, addressing data scarcity — a persistent bottleneck in training AI models for low-resource domains. In this setting, novelty scoring ensures varied outputs, while domain-specific constraints help maintain data validity.
Ultimately, our approach represents a step toward AI systems that do not just generate text but truly engage in creative synthesis. We hope to inspire further research in structured creativity and AI, highlighting the potential of structured recombination in augmenting human creativity and expanding the frontiers of AI-assisted innovation.
Appendix A Fine-Tuned Sentence-BERT Model
Our initial experiments with the standard sentence-level Sentence-BERT (SBERT) model revealed that it tends to group recipes emphasizing textual instructions while overlooking ingredients. As a result, it fails to distinguish broad categories (e.g., salad vs. soup) and struggles with finer-grained distinctions (e.g., carrot cake vs. cheesecake).
To better handle recipe similarity, we fine-tuned a Sentence-BERT model555all-distilroberta-v1. . Our fine-tuning dataset consisted of 30K pairs of recipes with their new similarity scores, which equally weighted the original Sentence-BERT score and a Ruzicka-based similarity of the two recipes’ ingredient lists as computed by Mizrahi and Shahaf (2021). The dataset was divided into three equal subsets: (1) 10K pairs of recipes, each pair representing instances of the same dish, (2) 10K pairs of recipes from different dishes within the same category (e.g., carrot cake and cheesecake), and (3) 10K pairs of recipes from entirely different categories (e.g., a dessert and a salad).
To preserve the original model’s capabilities and avoid overfitting, we fine-tuned it for one epoch, reserving 5% of the pairs for validation, achieving a validation accuracy of 92–95%. Following Sentence-BERT fine-tuning guidelines, we employed 10% warm-up steps, used cosine similarity loss as our loss function, and set the maximum sequence length to 512 tokens.
This fine-tuning procedure allowed us to obtain more accurate recipe embeddings that account for both textual instructions and ingredient overlap. Figure 9 presents a cosine similarity heatmap of 40 recipes before and after fine-tuning: 10 highly similar recipes for Carrot Cake, 10 for Truffles (desserts), 10 for Dumplings, and 10 for Pizza (main dishes). As shown, the original Sentence-BERT model exhibits high similarity scores across all dish pairs, including an unexpectedly high similarity between Carrot Cake and Pizza recipes. In contrast, the fine-tuned model retains high similarity scores within the same dish while improving differentiation between dishes within the same category and across categories.
Figure 9: Cosine similarity heatmaps of 40 recipes before and after fine-tuning: (A) Sentence-BERT before fine-tuning, (B) Sentence-BERT after fine-tuning. The original Sentence-BERT model shows uniformly high similarity across all dish pairs, while the fine-tuned model maintains high similarity within the same dish and improves differentiation across categories.
Appendix B Text to Tree Parser Details
Parsing the Ingredients.
In the first subtask, we instructed the model to parse the ingredients. For each ingredient line, the model extracted (1) the ingredient name, (2) whether the ingredient contributes to the dish’s structural core (e.g., lasagna sheets in lasagna) or its flavor (e.g., lemon in lemon pie), and (3) a simplified base form of the ingredient (e.g., “basil” → “herb,” “walnut” → “nut”). To evaluate accuracy, we sampled 200 random ingredient lines, and one of us checked whether the parsed information matched the ground truth. We obtained 95% accuracy for ingredient name parsing, 97.5% for the structural vs. flavor reference, and 96% for ingredient base form conversion. These results suggest that GPT-4o performs reliably in parsing ingredients.
Simplifying the Instructions.
Next, we asked the model to simplify each sentence in a recipe’s instructions while preserving essential content. Specifically, we instructed it to remove details about quantities, sizes, and descriptive elements, ensuring that each simplified sentence contained exactly one action (placed at the beginning) and that ambiguous instructions were converted into active forms (e.g., “bring to a boil” → “boil”). For example, the instruction “Sprinkle salt over the basil and mozzarella and return to the broiler for 1 to 2 minutes, until the cheese is melted and bubbling.” was simplified into “Sprinkle salt over basil, mozzarella. Broil until cheese melts”. We tested this phase on 50 sampled recipes, resulting in 451 instruction sentences. One of us reviewed these sentences and found that 423 (93.79%) were simplified correctly. Common errors included using non-informative verbs, omitting a verb or ingredient, or accidentally merging two actions into a single sentence.
Translating the Recipe into a Tree.
Finally, we used GPT-4o’s coding proficiency to produce a directed tree representation in DOT, a graph description language. We provided the model with a one-shot example that included a simplified recipe text (dish name, a short list of ingredients, and simplified instructions) along with its corresponding DOT code, annotated with comments to guide the model in structuring the tree representation. To refine the output, we implemented a correction step, automatically removing problematic edges and instructing the model to reconsider them.
Appendix C Text to Tree Parser Prompts
This section includes the prompts we used to prompt GPT-4o to parse a recipe text into a tree representation. The system message for all prompts presented here was: “You are a cooking recipe parser”.
Parsing the Ingredients Prompt:
Given a recipe title, id, and ingredients, for each ingredient, determine: (1) Abbreviation: The shortest clear description. (2) Reference Type: Identify if the ingredient is for structure (‘structure’) or taste (‘taste’) of the dish. Ingredients impacting both are labeled as ’taste’. (3) Core Ingredient: Boolean indicating if the ingredient is essential to the identity of the dish (e.g., True for chocolate in chocolate cake). (4) Abstraction: Simplify the ingredient to its base form (e.g., ‘basil’ to ‘herb’, ‘walnuts’ to ‘nut’, ‘eggs’ to ‘egg’). Please return the results in the following JSON format only: {“recipe_id”: [(abbreviation, ref, core, abstraction), …], …}. INPUT: recipe_title, recipe_id, ingr_list, recipe_title, recipe_id, ingr_list, … OUTPUT:
Simplifying the Instructions Prompt:
Given the following cooking instructions, please simplify and shorten them as much as possible. Remove quantities, sizes, and descriptions. Ensure each verb initiates a new sentence, and that a sentence does not contain two verbs. Convert permissive or ambiguous instructions into active forms (e.g., “let cool” -> “cool”, “alternate layers” -> “layer”). Return output in JSON format with the key as ‘recipe_id’ and the value as the full simplified text. INPUT: {recipe_id: <instruction text>, …} OUTPUT:
Translating the Recipe into a Tree Prompt:
Title: … Ingredients: … Directions: … Code: … # end of code (the 1-shot example)
Title: <dish_name> Ingredients: <ingrendient_abbreviation_list> Directions: [i1] <1st_direction> [i2]<2nd_direction> … Code:
Tree Correction Prompt:
You are provided with the title, ingredients, and directions of a recipe, along with a partial Dot code that represents the recipe’s tree structure. The Dot code is missing some edges. Additionally, you will receive names of nodes for which these connections are missing. For each provided node name, add exactly one edge from this node to the action node that uses it (if it is an ingredient) or processes its outcome (if it is an action). Please return only the Dot code for these specific edges, including necessary comments, and exclude any additional text. Title: <dish_name> Ingredients: <ingredient_abbreviation_list> Directions: [i1] <1st_direction> [i2] <2nd_direction> … Partial Dot code: <dot_code> Name of nodes with missing edges: <node_names> OUTPUT:
Appendix D Tree Edit Distance Implementation
In this appendix, we describe our approach for computing the minimal edit distance between recipe trees. As mentioned in Section 4.3, we employ the Zhang–Shasha algorithm Zhang and Shasha (1989), which extends the well-known string edit distance approach to ordered labeled trees. Specifically, a labeled tree is one in which each node is assigned a symbol from a fixed finite alphabet, and an ordered tree is one in which each set of siblings has a defined left-to-right order. While computing tree edit distance for unordered trees is NP-hard and even MAX SNP-hard Arora et al. (1992), the Zhang–Shasha algorithm provides a polynomial-time solution for the ordered case. To make our labeled recipe trees compatible with this approach, we impose an ordering on sibling nodes by sorting them lexicographically according to their labels.
We further adjust edit costs to encourage matching analogous nodes across recipes. Specifically, we allow a node to be substituted by another only if both nodes share the same type (i.e., both are ingredient nodes or both are action nodes). Moreover, a substitution between two nodes incurs zero cost if they have the same label, or a small fixed cost if their labels share the same abstract meaning. For example, two ingredients both categorized as “herb” may replace one another at a lower cost than an “herb” and a “liquid” ingredients. Similarly, two action nodes categorized under “heat application” are more likely to substitute for each other than an action node categorized under “heat application” and another one that is categorized under “flavor enhancement”.
To determine whether two ingredient nodes share the same abstraction, we use the ingredient abstraction obtained from parsing the ingredients (see Appendix B). To determine whether two action nodes share the same abstraction (e.g., “heat application” for “bake” and “microwave”), we collected the 250 most common action verbs in recipes and created a hierarchy, grouping these verbs into categories such as heat application, preparation, positioning, flavor enhancement, etc.
Formally, let $T_{1}$ and $T_{2}$ be two recipe trees composed of ingredient and action nodes. We allow the operations of insertion, deletion, and update. The cost of insertion or deletion is set to 100, whereas the update cost depends on whether the two nodes share the same type and the same label or abstraction. If they have the same type and the same label, the cost is 0; if they share the same type but only the same abstraction, it is 5; otherwise, the cost is $\infty$, indicating an infeasible substitution. This cost scheme encourages the edit distance algorithm to favor substituting analogous parts over insertions and deletions, resulting in more semantically meaningful transformations.
As noted in Section 4.3, stopping at different points in the transformation process can create unique dishes (see Figure 4). Additionally, shuffling the order of edits can generate entirely new intermediate ideas. To produce more coherent results, we impose a partial order on the shuffled operations. We prioritize inserting and updating key flavor ingredients from the target recipe (e.g., “lemon” in a lemon pie) so they appear earlier in the transformation. At the same time, we delay deleting or updating structural ingredients from the source recipe (e.g., “lasagna sheets” in lasagna) to preserve its core structure. We determine which ingredients contribute to flavor and which to structure during the parsing phase (see Appendix B). This approach helps maintain the essential characteristics of both dishes, integrating distinct flavor components from the target while retaining the structural integrity of the source. Notably, reversing the transformation (dish B → dish A) results in a different edit sequence, leading to distinct new recipes.
To ensure the dishes are indeed recombinations, we discard any recipe that lacks at least one essential ingredient from both original dishes (if such an ingredient exists). We define an essential ingredient as one that appears frequently in recipes for a given dish but is not broadly common across all recipes (e.g., lasagna sheets in lasagna). Additionally, we remove ideas that are too similar to the seed recipes. To ensure that the combined tree preserves cross-dish inspiration, we require that at least 30% of its elements (nodes and edges) come from each original recipe.
Appendix E Identifying Conflicting Ingredients
We remove ingredients that appear to collide in their flavors. Specifically, we look for pairs of ingredients that seldom appear together in the recipe repository, treating their pairing as uncommon and attempt to determine whether it might be a creative success or a failure.
To do so, we rely on two external datasets. First, we use flavorDB Garg et al. (2018), which catalogs taste molecules for a wide range of raw ingredients such as fruits, vegetables, and fish. Inspired by this dataset owner’s claim that two ingredients pair well if they share a larger proportion of taste molecules, we define a Jaccard-based pairing score between raw ingredients. Since flavorDB does not cover processed ingredients (e.g., lasagna sheets that consist of flour, water, and eggs), we also use FoodData Central Fukagawa et al. (2022) to infer their raw components. We define the pairing score between two composite ingredients as the low- est score among their constituent raw-ingredient pairs. After exploration, we chose 0.3 as our value threshold. If the score falls below 0.3, we consider the pairing problematic.
Appendix F Tree to Text Prompts
This section includes the prompts we used to prompt GPT-4o to translate structured tree representation back into natural language. The system message for all prompts presented here was: “‘You are a cooking expert”.
Translate Tree into Raw Recipe Prompt:
Given the following DOT code, which represents a recipe graphically by defining ingredient nodes, action nodes, and their interconnections, translate the structure into a natural language recipe. The DOT code maps each ingredient to specific actions, and it outlines the order of these actions to demonstrate the cooking process. DOT CODE: “‘<recipe_idea_dot_code>”’ Convert this structured representation into a detailed cooking recipe in natural language. Requirements: (1) Output should only include the title, ingredients with quantities, and sequential instructions. (2) Avoid any explanatory comments or embellishments. OUTPUT:
Find Issues and Correct Recipe Prompts:
Step I: Review the recipe provided below, which is written in natural language. Identify and list any potential issues with it, excluding any concerns related to unconventional ingredient combinations. Please provide only a list of potential issues without revising the recipe. RECIPE: “‘<GENERATED RECIPE>”’
Step II: Please edit the recipe to address the identified issues. Ensure the recipe remains as a single, unified component. Output only the corrected version of the recipe. OUTPUT:
Summarize Recipe Prompt:
Please summarize the following recipe in a few sentences: (1) Start with a super concise description of the dish, focusing *only* on its final result. (2) Then, provide a summary of the recipe, including its main components, actions, and all the ingredients used. Use a descriptive tone for this part, avoiding imperative sentences. RECIPE: “‘<full_recipe>”’
Review Ingredients Prompt:
You are given a description of a creative recipe. CREATIVE RECIPE DESCRIPTION: “‘<creative_recipe_description>”’ Your task is to preserve the creative ingredients in the recipe while suggesting the removal or substitution of ingredients that might negatively impact the dish’s flavor. You should: (1) Recognize the unique and unusual ingredients that contribute to the creativity of the dish. (2) Systematically compare all pairs of ingredients in the dish and identify ingredients that have a clear, strong clash with each other due to conflicting flavors. Be thorough and ensure that you include all possible pairs of ingredients that have a strong clash. (3) Based on the identified strong clashes, suggest removals and substitutions of ingredients to avoid clashes, while preserving the creative aspects of the dish. Return only the following JSON output format: {“dish_ingredients”: <list of strings: the full list of ingredients in the dish>, “creative_ingrs”: <list of strings: the list of ingredients that contribute creatively to the dish>, “flavor_clashes”: <list of string pairs: the clashing ingredients>, “removals”: <list of strings: the list of ingredients to remove>, “substitutions”: <list of string pairs: ingredients to substitute - (ingr1, ingr2) means ‘replace ingr1 in ingr2’>}
Increase Readability Prompt:
Given the following recipe: (1) Remove the following ingredients: <bad_ingredients>. (2) Make the following ingredient substitutions: <required_substitutions>. (3) Split its ingredients and instructions into distinct sections to improve readability (e.g., “mix dry ingredients”, “assemble”, etc.). You can change the order of lines but keep the content unchanged. “‘<full_recipe>”’
Appendix G Chosen Prompts for Experiments
In this section, we present the prompts used to guide GPT-4o in generating recipes for both experiments.
Experiment 1 Prompt:
You are a chef at a fusion restaurant that excels in creating delightful and unexpected combinations of classic dishes. Your task today is to design an innovative recipe that merges the intricate layers of {dish1} with the rich decadence of {dish2}. Develop a comprehensive recipe that includes: (1) A unique name that embodies the essence of this fusion dish. (2) A detailed list of ingredients. (3) Step-by-step cooking and assembly instructions, highlighting inventive cooking techniques or unusual ingredient interactions. Promote bold experimentation with flavors and textures to create a dish that is both surprising and satisfying.
Following prompt: Design another different innovative recipe that merges the intricate layers of {dish1} with the rich decadence of {dish2}.
Experiment 2 Prompt:
What is the most creative and out-of-the-box recipe you can create?
Following prompt: Create a fresh, unique recipe that differs from the previous ones but matches their level of creativity.
Appendix H Human Experiment Questionnaire
  1. 1.
    Do the instructions in this recipe make sense? A recipe that doesn’t make a lot of sense contains technical issues. Issues could be minor (for example, mixing an already-mixed salad) or major (not cooking raw chicken). {scale (1-5): 1: Throw away the recipe, too many changes needed, 2: Need to change most of the recipe, 3: Requires a lot of changes, 4: Almost perfect, requires some minor changes, 5: Makes perfect sense, could cook it as is}
  2. 2.
    Do the combination of ingredients in this recipe make sense? {scale: same}
  3. 3.
    How similar is this recipe to others you have seen or used? {scale (1-5): 1: Very similar to many recipes, 2: Similar but not very common, 3: Somewhat similar to others, 4: Quite different from most recipes, 5: Very different, highly unusual}
  4. 4.
    How novel is the way the instructions are combined in this recipe compared to typical recipes? {scale: same}
  5. 5.
    How novel is the combination of the ingredients in this recipe compared to typical recipes? {scale: same}
  6. 6.
    Assuming a cook followed this recipe, would people want to taste it? {scale (1-5): 1: Never, 2: Only if they have to, 3: Only if they’re really hungry, 4: They’d probably try it, 5: Yes, definitely}
  7. 7.
    Assuming a cook followed this recipe after the required modifications, would people want to taste it? {scale: same}
  8. 8.
    How original do you find this recipe overall? (after the required modifications) {scale (1-5): 1: Not original at all, 2: Slightly original, 3: Somewhat original, 4: Fairly original, 5: Extremely original and creative}